Evaluate the limits

Lim

X to 2

of

Xcube -8 over x-2


1
Expert's answer
2021-05-19T18:06:26-0400

We need to calculate "\overset{lim}{(x\to2)}" "\frac {(x^3-8)} {(x-2)}"

So here this is in "\frac 0 0" form after putting 2 in place of x. So L-Hospital's rule is applicable here.

= "\overset {lim}{(x\to2)}" "\frac {\frac {d} {dx} (x^3-8)} {\frac {d} {dx} (x-2)}"


="\overset {lim} {x\to2} { ( \frac {3x^2} {1})}" (As "\frac{d} {dx} (x^n)=n.x^{n-1} )" & "(\frac {d} {dx} (constant)=0)"


= "\frac {3(2)^2} {1}"

="3*4=12" (Ans)